Estimate the work necessary to compress the air in an air-compressor cylinder from a pressure of 100 kPa to 2000 kPa. The initial volume is 1000cm^3. An isothermal process is to be assumed.
Answer #1

PV = RT ====> RT = PV = (100000*1000*10^-6) = 100

Now work done = RT*ln(P1/P2) = 100*ln(100/2000) = -300 J (-ve sign for compression)